STATE-TO-STATE ROTATIONAL AND VIBRATIONAL ENERGY TRANSFERS FOLLOWING VIBRA-TIONAL EXCITATION OF (1010^{0}0^{0}) AND (0112^{0}0^{0}) IN THE GROUND ELECTRONIC STATE OF ACETY-LENE

\underline{JIANDE HAN}, KEITH FREEL, and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

We have examined state-to-state rotational and vibrational energy transfers for the vibrational levels (1010^{0}0^{0}) and (0112^{0}0^{0}) of C_{2}H_{2} in the ground electronic state at ambient temperature. Measurements were made using a pulsed IR - UV double resonance technique. Total removal rate constants and state-to-state rotational energy transfer rate constants have been characterized for certain even-numbered rotational levels from J = 0 to 12 within the two vibrational modes. The measured state-to-state rotational energy transfer rate constants were fit to some energy-based empirical scaling and fitting laws, and the rate constants were found to be best reproduced by the statistical power-exponential gap law (PEGL). The measured rate constants were then further evaluated by a kinetic model which simulated the experimental spectra by solving simultaneous first order differential rate equations. Some rotationally-resolved vibrational energy transfer channels were also observed following excitation of (1010^{0}0^{0}). The vibrational relaxation channels were found to contribute less than 30% to the total removal rate constants of the measured rotational levels for both of the studied vibrational states.